# Asteroseismic Structure Inversions of Main-Sequence Solar-like Oscillators with Convective Cores

### v1

@@ 1 Introduction
In Figure 1, we show the existing sample of main-sequence solar-like oscillators studied using asteroseismic structure inversions. Buchele et al. (2024), henceforth B24, presented results for 12 stars with radiative cores, including the solar analogues 16 Cyg A and B which were also studied by Bellinger et al. (2017) and Buldgen et al. (2022b). Structure inversions have also been used to study a main-sequence star with a small convective core (Bellinger et al., 2019) and two stars evolved enough to exhibit mixed modes (Kosovichev & Kitiashvili, 2020). All three of these stars are in the sample presented here, where we extend the work of B24 to cover main-sequence solar-like oscillators with convective cores observed by Kepler.
Figure 1: Hertzsprung–Russell diagram for main-sequence solar-like oscillators, with inversion results available. Stars have been categorized based on their inversion results in this work and in B24, represented by the color of the symbol. Stars with other inversion results available are indicated with larger open symbols. The uncertainties of the luminosity values are smaller than the points. Stellar evolutionary tracks of several masses are shown for reference.
@@ 2 Forward Modeling
As structure inversions infer differences between a star and a model, the first step is to find a suitably close reference model, typically the best-fit model from a grid-based modeling or optimization procedure. The process of finding such a model is called forward modeling. To find our reference models, we used a grid-based method similar to that used by B24. We constructed a grid of 24,530 tracks using the stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The details of the grid that are unchanged between this work and B24 are detailed in Appendix A.1. There are two important changes which we discuss here. It is well known that including diffusion and gravitational settling of elements without also including the effects of radiative levitation produces models with unrealistic surface abundances in this mass range (e.g. Michaud et al., 2015; Deal et al., 2018). However, including radiative levitation significantly increases the computation time of models, such that it would be difficult to compute the number of models necessary to cover the parameters space of the observations. We chose to compromise and evolve our tracks without including diffusion, settling, or radiative levitation. Additionally, since we are dealing with stars with convective cores, we use the exponential overshooting scheme implemented in MESA where the overshoot region is treated as fully mixed without changing the thermal gradient. This is described in detail in Paxton et al. (2011). The parameters varied in this grid are mass, initial helium abundance, initial metallicity, mixing length parameter, and overshooting parameter. To cover the parameter space efficiently, we varied each parameter using a Sobol sequence (see Appendix B of Bellinger et al., 2016; Sobol’, 1967) within the ranges listed in Table 1. For each model in the grid, we calculated the adiabatic frequencies using GYRE (Townsend & Teitler, 2013). We then scanned the grid to find the model parameters that best fit the frequencies, effective temperature, and metallicity by minimizing:
@@
While scanning the grid, we interpolated along each track, but not between tracks. This is the same method as that of B24 with one change — in this work we interpolated in age instead of central hydrogen abundance, as the central hydrogen abundance does not decrease monotonically in stars where a convective core emerges after the zero-age main-sequence. From this procedure, we obtain the best-fit parameters, which are then used to calculate the reference model of the structure inversions. Our values of $T_{\rm{eff}}$ and [Fe/H] come from Furlan et al. (2018); Mathur et al. (2017); Morel et al. (2021), with the specific source for each star given in Appendix A.2. We also provide, in Appendix A.2, the parameters of our best-fit model as well as a comparison to the parameters reported in Silva Aguirre et al. (2017). Additionally, we provide both the FGONG structure file and the inlist used to generate each model at link to be added upon publication.
@@ 3 Structure Inversions
where $R$ and $M$ are the stellar radius and mass, respectively. This choice of variables is well suited for asteroseismic targets (Basu, 2003; Bellinger et al., 2020), as the oscillations are mostly insensitive to $Y$. This makes it easier to isolate the differences due to a change in $\hat{u}$.
@@
Such an equation can be written for each mode $i$, where the index $i$ of the mode again corresponds to a specific pair of $n,l$. The relative frequency difference ($\delta\hat{\nu}_{i}/\hat{\nu}_{i}$) is related to the structure differences between the model and the observed star through the mode kernel functions $K_{i}$. These mode kernels are known functions of the reference model, found though a linear perturbation of the oscillation equations (for more details, see Gough & Thompson 1991, Kosovichev 1999, or Thompson & Christensen-Dalsgaard 2002). As we are inverting for dimensionless structure variables we need to use the difference in dimensionless frequency ($\hat{\nu}$). These differences are calculated by scaling the dimensional frequency differences by the large frequency separation, $\Delta\nu$:
@@
This works because $\Delta\nu$ carries the same dependence on the stellar mass and radius as the frequencies. The full derivation of Equation 6 can be found in Appendix B2 of B24. We note that the model frequencies $\nu_{i,\rm{mod}}$ in this expression have been corrected for surface effects using the two-term formulation of Ball & Gizon (2014). We also calculate $\Delta\nu_{\rm{mod}}$ after applying the surface term correction.
We choose to correct for dimensional differences and surface effects when calculating the frequency differences. These effects, however, can be handled in different ways. Alternative approaches to handling dimensional differences include using a different correction method (Bellinger et al., 2021), including the mean density in the fitting procedure (Buldgen et al., 2022b), or adding a term to Equation 5 (Kosovichev & Kitiashvili, 2020). Following the arguments outlined in Appendix B2 of B24, we expect that the inversion procedure will suppress the effects of differences in mean density regardless of the correction method used. Bellinger et al. (2019) shows this explicitly as inversions using models of different masses and radii return the same results. The surface term can also be accounted for during the inversion by adding a term to Equation 5 (e.g., Equation 3 of Bellinger et al., 2016). We have tested this approach and found no difference in the final inversion results.
@@
$\sum_{i}^{N}c_{i}\frac{\delta\nu_{i}}{\nu_{i}}=\int\mathcal{K}_{r_{0}}\frac{\delta\hat{u}}{\hat{u}}\,\textrm{d}r+\int\mathcal{C}_{r_{0}}\,\delta Y\,\textrm{d}r.$ (7)
Here $c_{i}$ are known as the inversion coefficients, $\mathcal{K}=\sum_{i}^{N}c_{i}K_{i}^{(\hat{u},Y)}$ is called the averaging kernel, and $\mathcal{C}=\sum_{i}^{N}c_{i}K_{i}^{(Y,\hat{u})}$ is the cross-term kernel. When the inversion coefficients are chosen such that $\mathcal{K}$ is localized around $r_{0}$ and normalized to 1, and the amplitude of $\mathcal{C}$ is small everywhere, then Equation 7 reduces to
$\sum_{i}^{N}c_{i}\frac{\delta\nu_{i}}{\nu_{i}}\approx\int\mathcal{K}_{r_{0}}\frac{\delta\hat{u}}{\hat{u}}\>\textrm{d}r\approx\left<\frac{\delta\hat{u}}{\hat{u}}\right>_{r_{0}}.$ (8)
@@
To find the inversion coefficients, we used the method of multiplicative optimally localized averages (MOLA), which constructs the averaging kernel by penalizing any amplitude away from the target radius. For details on the implementation of MOLA, see Basu & Chaplin (2017, Chapter 10). In this process, we must choose two trade-off parameters: $\beta$, the cross-term suppression parameter, and $\mu$, the error suppression parameter. We chose our parameter values using the same method as B24. Briefly, this method sets $\beta=0$ as the choice of $Y$ as the second variable naturally suppresses the amplitude of the cross-term kernel. We then chose a value of $\mu$ that could correctly recover the known values of $\delta\hat{u}/\hat{u}$ between our reference model and a small set of calibration models. These models came from our grid and have slightly larger values of $\chi^{2}_{\rm{fit}}$ than our reference model.
@@ 4 Results
Figure 2: Inversion results for one star in each category. Each shows the relative differences in $\hat{u}$ between observation and best-fit model inferred by the inversions, in the sense of (Star -- Model)/Model. The vertical error bars indicate the uncertainty of each inversion result from the propagation of the uncertainty of the observed frequencies. The horizontal error bars represent the full width at half maximum of the averaging kernel. The dashed horizontal line indicates complete agreement between the model and observations; points above this line imply that $\hat{u}$ of the star is larger than that of the model. The color bar indicates the statistical significance of the inferred difference, with lighter colors showing more significant results. We also provide the mass ($M$, in $M_{\odot}$), initial helium mass fraction ($Y_{\rm{init}}$), initial metallicity ($Z_{\rm{init}}$), central hydrogen mass fraction ($X_{c}$), and mass of the convective fore ($M_{cc}$, in $M_{\odot}$) of each model.
Around half (24) of the stars fall into category (A). These models still show significant differences in the oscillation frequencies, even after correcting for the surface term, which suggests that the structure differences are either smaller than the observational uncertainties at the resolution given by the structure inversions or that the structure differences are at a location unable to be probed by the inversions. Of the stars showing significant disagreement, 11 are in category (H), 6 are in category (L), and 1 each in categories (LH) and (HL). Using the $\chi^{2}_{\rm{inv}}$ parameter defined in B24, we search for correlations with a variety of model parameters and observations. In contrast to the earlier work, we find no significant correlations. One problem with the $\chi^{2}_{\rm{inv}}$ metric is that it only measures the significance of the inversion results, not whether the differences inferred are positive or negative. To account for this, we also look for correlations between the model parameters and $a\chi^{2}_{\rm{inv}}$ where $a=-1(+1)$ for stars where the most significant inferred difference is negative (positive). We also find no significant correlations in this case.
@@
Figure 3: Information about the best-fit model of KIC 6225718. The top left plot shows the frequency échelle diagram comparing the frequencies of the reference model to the observations after correcting for surface effects. The top right plot shows the frequency separation ratios of the reference model and observations. The ratios of the reference model have been computed from the uncorrected model frequencies. The bottom left plot shows the averaging kernels and the bottom right the cross-term kernels. Note that the $y$-axis scale differs by an order of magnitude between the two plots on the bottom row.
@@ 4.1 KIC 6225718
The first star we discuss in detail is KIC 6225718. Our inversion results for this star are shown in Figure 2. We show in Figure 3 the frequency échelle diagram and frequency separation ratios of our best-fit model, as well as, our averaging and cross-term kernels. This star has already been studied using structure inversions by Bellinger et al. (2019) which allows us to compare our results. Structure inversions infer differences relative to a given reference model using a given set of averaging kernels. As such, any comparison of different inversion results must be considered in the context of the reference model and averaging kernels used. The averaging kernels we use are very similar to those used in Bellinger et al. (2019). In comparing the structure of our reference model to the original model used by Bellinger et al. (2019), we find that ours has a slightly larger convective core (with our convective boundary located at $r/R\approx 0.08$ instead of $r/R\approx 0.04$). We attribute this difference to our inclusion of core overshoot. Bellinger et al. (2019) also constructed a model with core overshoot which resulted in a convective core boundary closer to our reference model. The inversion results for this model are not reported, however, the authors note that the inversion results were the same as their original model. Thus we conclude that our inversion results can be directly compared to those in Bellinger et al. (2019).
Both works find that the model $\hat{u}$ is too low in the outermost regions probed by inversions and too high in the innermost regions, with the crossover occurring around $r/R\approx 0.25$. Bellinger et al. (2019) find a maximum difference at $r/R\approx 0.1$, while our maximum difference is found around $r/R\approx 0.05$. However, the higher uncertainties at these target radii mean that this difference in inversion results is not significant. Thus, despite slight differences in the fitting and inversion procedures, we find the same disagreement in internal structure.
@@ 4.2 KIC 10162436 and KIC 5773345
Kosovichev & Kitiashvili (2020) present inversion results for two stars that are also in our sample: KIC 10162436 and KIC 5773345. Our results for these stars are shown in Figure 4. Directly comparing our inversion results to theirs is slightly more difficult than with Bellinger et al. (2019). For both stars, they find mixed modes with $l=1$ and $l=2$. Our model of KIC 10162436 has one mixed $l=2$ mode, which we exclude from our inversions, and no $l=1$ mixed mode. In the case of KIC 5773345, our model shows no mixed modes at all. In addition, the frequency differences, even of pure acoustic modes, between their models and the observations are significantly larger than ours. These differences in frequencies and mode character, which we attribute to differences in the modeling procedure [^1] Kosovichev & Kitiashvili (2020) use the parameters from the YMCM modeling pipeline presented in Silva Aguirre et al. (2017), including the mixing length parameter and stellar age, to compute a model using MESA. However, the default nuclear reaction rates and formulation of mixing length theory implemented in MESA differ from those used in the YMCM models. It is unclear if the MESA defaults were changed to match the physics used in the YMCM models. If these changes were not made the YMCM parameters may not result in a best-fit model., suggest that the structure of the reference models used in Kosovichev & Kitiashvili (2020) and this work are different. Additionally, while Kosovichev & Kitiashvili (2020) do not show their averaging kernels, the spread indicated by their horizontal error bars are much wider than ours. Thus we cannot directly compare our inversion results to those given in Kosovichev & Kitiashvili (2020). Nevertheless, we note that we infer differences in $\hat{u}$ of similar magnitude for both KIC 10162436 and KIC 5773345.
@@
Figure 4: Frequency differences, frequency separation ratios, and inversion results for our models of the two stars shown in Kosovichev & Kitiashvili (2020). The top row shows the results for KIC 10162436. The figure on the left shows the relative frequency differences, after correcting for surface effects. The center panel shows the frequency separation ratios of the observed star and our reference model (computed using uncorrected model frequencies). The right panel shows the inversion results, where all symbols and colors have the same meaning as in Figure 2. The lower row shows the same information for KIC 5773345.

### v2

@@ 1 Introduction
In Figure 1, we show the existing sample of main-sequence solar-like oscillators studied using asteroseismic structure inversions. Buchele et al. (2024), henceforth B24, presented results for 12 stars with radiative cores, including the solar analogues 16 Cyg A and B which were also studied by Bellinger et al. (2017) and Buldgen et al. (2022b). Structure inversions have also been used to study a main-sequence star with a small convective core (Bellinger et al., 2019) and two stars evolved enough to exhibit mixed modes (Kosovichev & Kitiashvili, 2020). All three of these stars are in the sample presented here, where we extend the work of B24 to cover main-sequence solar-like oscillators with convective cores observed by Kepler. Our primary goals of this work are to determine which stars have observational data suitable for inversions and to examine how well models created using common modeling choices reproduce the structure of observed stars.
Figure 1: Hertzsprung–Russell diagram for main-sequence solar-like oscillators with inversion results available. Stars have been categorized based on their inversion results in this work and in B24, represented by the color of the symbol. Stars with other inversion results available are indicated with larger open symbols. The uncertainties of the luminosity values are smaller than the points. Stellar evolutionary tracks of several masses are shown for reference.
@@ 2 Forward Modeling
As structure inversions infer differences between a star and a model, the first step is to find a suitably close reference model, typically the best-fit model from a grid-based modeling or optimization procedure. The process of finding such a model is called forward modeling. To find our reference models, we used a grid-based method similar to that used by B24. We constructed a grid of 24,530 tracks using the stellar evolution code MESA (Paxton et al., 2011, 2013, 2015, 2018, 2019; Jermyn et al., 2023). The details of the grid that are unchanged between this work and B24 are provided in Appendix A.1. There are two important changes which we discuss here. It is well known that including diffusion and gravitational settling of elements without also including the effects of radiative levitation produces models with unrealistic surface abundances in this mass range (e.g. Michaud et al., 2015; Deal et al., 2018). However, including radiative levitation significantly increases the computation time of models, such that it would be difficult to compute the number of models necessary to cover the parameters space of the observations. We chose to compromise and evolve our tracks without including diffusion, settling, or radiative levitation. Additionally, since we are dealing with stars with convective cores, we use the exponential overshooting (Freytag et al., 1996; Herwig, 2000) scheme implemented in MESA where the overshoot region is treated as fully mixed without changing the thermal gradient. This is described in detail in Paxton et al. (2011). The parameters varied in this grid are mass, initial helium abundance, initial metallicity, mixing length parameter, and overshooting parameter. To cover the parameter space efficiently, we varied each parameter using a Sobol sequence (see Appendix B of Bellinger et al., 2016; Sobol’, 1967) within the ranges listed in Table 1. For each model in the grid, we calculated the adiabatic frequencies using GYRE (Townsend & Teitler, 2013). We then scanned the grid to find the model parameters that best fit the frequencies, effective temperature, and metallicity by minimizing:
@@
While scanning the grid, we interpolated along each track, but not between tracks. This is the same method as that of B24 with one change — in this work we interpolated in age instead of central hydrogen abundance, as the central hydrogen abundance does not decrease monotonically in stars where a convective core emerges after the zero-age main-sequence. From this procedure, we obtain the best-fit parameters, which are then used to calculate the reference model of the structure inversions. Our values of $T_{\rm{eff}}$ and [Fe/H] come from Furlan et al. (2018); Mathur et al. (2017); Morel et al. (2021), with the specific source for each star given in Appendix A.2. We also provide, in Appendix A.2, the parameters of our best-fit model, a comparison to the parameters reported in Silva Aguirre et al. (2017), and a comparison of our $\chi^{2}_{\rm{fit}}$ distribution to that obtained using the YMCM pipeline of Silva Aguirre et al. (2017). Additionally, we provide both the FGONG structure file and the inlist used to generate each model at https://zenodo.org/records/15341350.
@@ 3 Structure Inversions
where $R$ and $M$ are the stellar radius and mass, respectively. This choice of variables is well suited for asteroseismic targets (Gough & Kosovichev, 1993; Basu, 2003; Bellinger et al., 2020), as the oscillations are mostly insensitive to $Y$. This makes it easier to isolate the differences due to a change in $\hat{u}$.
@@
Such an equation can be written for each mode $i$, where the index $i$ of the mode again corresponds to a specific pair of $n,l$. The relative frequency difference ($\delta\hat{\nu}_{i}/\hat{\nu}_{i}$) is related to the structure differences between the model and the observed star through the mode kernel functions $K_{i}$. These mode kernels are known functions of the reference model, found through a linear perturbation of the oscillation equations (for more details, see Gough & Thompson 1991, Kosovichev 1999, or Thompson & Christensen-Dalsgaard 2002). There are two potential sources of frequency differences that are not accounted for in the mode kernels: surface effects and differences in mean density. We correct for the surface term during the calculation of our frequency differences using the same correction applied when finding our reference model, the two-term correction of Ball & Gizon (2014). We account for mean density differences by inverting for dimensionless structure variables. This requires us to use the difference in dimensionless frequency ($\hat{\nu}$). These differences are calculated by scaling the dimensional frequency differences by the large frequency separation, $\Delta\nu$:
@@
This works because $\Delta\nu$ carries the same dependence on the stellar mass and radius as the frequencies. The full derivation of Equation 6 can be found in Appendix B2 of B24. In this expression $\Delta\nu_{\rm{mod}}$ is calculated after applying the surface term correction. We note that both of these effects can be handled in different ways.
Alternative approaches to handling dimensional differences include using a different correction method (e.g. Roxburgh et al., 1998; Basu, 2003; Bellinger et al., 2021), including the mean density in the fitting procedure (e.g. Buldgen et al., 2022b), or adding a term to Equation 5 (e.g. Gough & Kosovichev, 1993; Kosovichev & Kitiashvili, 2020). Following the arguments outlined in Appendix B2 of B24, we expect that the inversion procedure will suppress the effects of differences in mean density regardless of the correction method used. Bellinger et al. (2019) shows this explicitly as inversions using models of different masses and radii return the same results. The surface term can be accounted for during the inversion by adding a term to Equation 5 (e.g. Rabello-Soares et al., 1998; Bellinger et al., 2016; Buldgen et al., 2022b). We have tested this approach and found no difference in the final inversion results.
@@
$\sum_{i}^{N}c_{i}\frac{\delta\hat{\nu}_{i}}{\hat{\nu}_{i}}=\int\mathcal{K}_{r_{0}}\frac{\delta\hat{u}}{\hat{u}}\,\textrm{d}r+\int\mathcal{C}_{r_{0}}\,\delta Y\,\textrm{d}r.$ (7)
Here $N$ is the total number of modes, $c_{i}$ are known as the inversion coefficients, $\mathcal{K}=\sum_{i}^{N}c_{i}K_{i}^{(\hat{u},Y)}$ is called the averaging kernel, and $\mathcal{C}=\sum_{i}^{N}c_{i}K_{i}^{(Y,\hat{u})}$ is the cross-term kernel. When the inversion coefficients are chosen such that $\mathcal{K}$ is localized around $r_{0}$ and normalized to 1, and the amplitude of $\mathcal{C}$ is small everywhere, then Equation 7 reduces to
$\sum_{i}^{N}c_{i}\frac{\delta\hat{\nu}_{i}}{\hat{\nu}_{i}}\approx\int\mathcal{K}_{r_{0}}\frac{\delta\hat{u}}{\hat{u}}\>\textrm{d}r\approx\left<\frac{\delta\hat{u}}{\hat{u}}\right>_{r_{0}}.$ (8)
@@
To find the inversion coefficients, we used the method of multiplicative optimally localized averages (MOLA), which constructs the averaging kernel by penalizing any amplitude away from the target radius. For details on the implementation of MOLA, see Basu & Chaplin (2017, Chapter 10). In this process, we must choose two trade-off parameters: $\beta$, the cross-term suppression parameter, and $\mu$, the error suppression parameter. We chose our parameter values using the same method as B24. Briefly, this method sets $\beta=0$ as the choice of $Y$ as the second variable naturally suppresses the amplitude of the cross-term kernel. We then chose a value of $\mu$ that correctly recovers the known values of $\delta\hat{u}/\hat{u}$ between our reference model and a small set of calibration models. These models came from our grid and have slightly larger values of $\chi^{2}_{\rm{fit}}$ than our reference model.
@@ 4 Results
Figure 2: Inversion results for one star in each category. Each shows the relative differences in $\hat{u}$ between observation and best-fit model inferred by the inversions, in the sense of (Star -- Model)/Model. The vertical error bars indicate the uncertainty of each inversion result from the propagation of the uncertainty of the observed frequencies. The horizontal error bars represent the FWHM of the averaging kernel. The dashed horizontal line indicates complete agreement between the model and observations; points above this line imply that $\hat{u}$ of the star is larger than that of the model. The color bar indicates the statistical significance of the inferred difference, with lighter colors showing more significant results. We also provide the mass ($M$, in $M_{\odot}$), initial helium mass fraction ($Y_{\rm{init}}$), initial metallicity ($Z_{\rm{init}}$), central hydrogen mass fraction ($X_{c}$), and mass of the convective fore ($M_{cc}$, in $M_{\odot}$) of each model.
Around half (24) of the stars fall into category (A). These models still show significant differences in the oscillation frequencies, even after correcting for the surface term, which suggests that the structure differences are either smaller than the observational uncertainties at the resolution given by the structure inversions or that the structure differences are at a location unable to be probed by the inversions. Of the stars showing significant disagreement, 11 are in category (H), 6 are in category (L), and 1 each is in categories (LH) and (HL). Using the $\chi^{2}_{\rm{inv}}$ parameter defined in B24, we search for correlations with a variety of model parameters and observations. In contrast to the earlier work, we find no significant correlations. One problem with the $\chi^{2}_{\rm{inv}}$ metric is that it only measures the significance of the inversion results, not whether the differences inferred are positive or negative. To account for this, we also look for correlations between the model parameters and $a\chi^{2}_{\rm{inv}}$ where $a=-1(+1)$ for stars where the most significant inferred difference is negative (positive). We also find no significant correlations in this case.
@@
Figure 3: Information about the best-fit model of KIC 6225718. The top left plot shows the frequency échelle diagram comparing the frequencies of the reference model to the observations after correcting for surface effects. The top right plot shows the frequency separation ratios of the reference model and observations. The bottom left plot shows the averaging kernels and the bottom right the cross-term kernels. Note that the $y$-axis scale differs by an order of magnitude between the two plots on the bottom row.
@@ 4.1 KIC 6225718
The first star we discuss in detail is KIC 6225718. Our inversion results for this star are shown in Figure 2. We show in Figure 3 the frequency échelle diagram and frequency separation ratios of our best-fit model, as well as, our averaging and cross-term kernels. This star has already been studied using structure inversions by Bellinger et al. (2019) which allows us to compare our results. Structure inversions infer differences relative to a given reference model using a given set of averaging kernels. As such, any comparison of different inversion results must be considered in the context of the reference model and averaging kernels used. The averaging kernels we use are very similar to those used in Bellinger et al. (2019). We focus on the structure for our comparisons. We have obtained the reference models used in Bellinger et al. (2019) to test the variation of the modeling physics and compare them to our reference model in Figure 4. In all cases, the models of Bellinger et al. (2019) are more similar to each other than to our reference model, although the model with physics closest to our choices (overshoot without diffusion) is the closest to our reference model. This suggests that the differences between our models are due to differences in the fitting procedure.
[…]
Figure 4: Structure differences between the reference model for KIC 6225718 used in this work and those used in Bellinger et al. (2019) (B19). Each line shows the relative difference in $\hat{u}$ between our reference model and the model from Bellinger et al. (2019) constructed using the modeling physics indicated in the legend. We indicate 0 with a dashed horizontal line and plot our inversion results for this star in black points
@@ 4.2 KIC 10162436 and KIC 5773345
Kosovichev & Kitiashvili (2020) present inversion results for two stars that are also in our sample: KIC 10162436 and KIC 5773345. Our results for these stars are shown in Figure 5. Both of these stars, as well as the star discussed in the next section, show glitches in their frequency separation ratios. These go beyond the change of slope expected due to core overshooting (Deheuvels et al., 2016). Previous works (Lebreton & Goupil, 2012; Deal et al., 2023) have shown that similar glitches can be induced in the frequency separation ratios of models by including large amounts (1-2 times the pressure scale hight) of convective penetration at the base of the convection zone, although this is considered physically unrealistic. The inability of models to reproduce these glitches in the ratios is an additional suggestion that something is missing in our stellar models.
@@
Directly comparing our inversion results to theirs is slightly more difficult than with Bellinger et al. (2019). For both stars, they find mixed modes with $l=1$ and $l=2$. Our model of KIC 10162436 has one mixed $l=2$ mode, which we exclude from our inversions, and no $l=1$ mixed mode. In the case of KIC 5773345, our model shows no mixed modes at all. In addition, the frequency differences, even of pure acoustic modes, between their models and the observations are significantly larger than ours. We attribute these differences to differences in the modeling procedure. Kosovichev & Kitiashvili (2020) use the parameters from the YMCM modeling pipeline presented in Silva Aguirre et al. (2017), including the mixing length parameter and stellar age, to compute a model using MESA. However, the YMCM models were computed using YREC, a different stellar evolution code, and care must be taken when using best fit parameters, especially mixing length parameter and stellar age, across different codes to ensure that the implemented physics matches as closely as possible. In particular, the nuclear reaction rates and formulation of mixing length theory differ between YREC and the defaults used in MESA. It is unclear whether the authors of Kosovichev & Kitiashvili (2020) made the necessary changes to MESA to match the original YREC configuration. These differences likely explain the large frequency differences that Kosovichev & Kitiashvili (2020) find between the observed and modeled frequencies, despite the model parameters being the same. Taken together, these differences suggest that the structure of the reference models used in Kosovichev & Kitiashvili (2020) and this work are different. Additionally, while Kosovichev & Kitiashvili (2020) do not show their averaging kernels, the spread indicated by their horizontal error bars are much wider than ours. Thus we cannot directly compare our inversion results to those given in Kosovichev & Kitiashvili (2020). Nevertheless, we note that we infer differences in $\hat{u}$ of similar magnitude for both KIC 10162436 and KIC 5773345.
@@
Figure 5: Frequency differences, frequency separation ratios, and inversion results for our models of the two stars shown in Kosovichev & Kitiashvili (2020). The top row shows the results for KIC 10162436. The figure on the left shows the relative frequency differences, after correcting for surface effects. The center panel shows the frequency separation ratios of the observed star and our reference model. The right panel shows the inversion results, where all symbols and colors have the same meaning as in Figure 2. The lower row shows the same information for KIC 5773345.